OpenStudy (anonymous):

if a=3, b=-1 and c=0 determine the value of (a) abc (b) (a+b)x2 (c) ab over c

OpenStudy (anonymous):

abc=3(-1)(0)=0 2(a+b)=2(3-1)=2(2)=4 ab/c=-3/0 not defined

OpenStudy (anonymous):

in grade 8, need to know what the rules are behind that sum...confused. test tomorrow :(

OpenStudy (anonymous):

never divide by 0,this is the rule for grade 8... adding signed numbers shouldn't be a problem,multiplying them as well..

OpenStudy (anonymous):

please explain what you mean by a signed number?

OpenStudy (anonymous):

a number having a + or - before it

OpenStudy (anonymous):

\[\sqrt{64x ^{2}}y ^{4}\] -please can you explain how to simplify this

OpenStudy (anonymous):

\[\sqrt[3]{27a ^{9}}y ^{3} \] please can you explain how to simply this

OpenStudy (anonymous):

\[\sqrt{64x ^{2}y ^{4}}=\sqrt{8^{2}x ^{2}(y ^{2})^{2}}=8xy ^{2}\]

OpenStudy (anonymous):

\[\sqrt[3]{27a ^{9}y ^{3}}=\sqrt[3]{3^{3}(a ^{3})^{3}y ^{3}}=3a ^{3}y\]

OpenStudy (anonymous):

subtract \[2x ^{2}y ^{4}\] from \[3x ^{2}-7x+3\]

OpenStudy (anonymous):

never mind i got the answer

OpenStudy (anonymous):

simplify \[-5x-9x\]

OpenStudy (anonymous):

you should try first...

OpenStudy (anonymous):

k but i realy suck at maths im failing it

OpenStudy (anonymous):

if im right it should be -14

OpenStudy (anonymous):

-14x

OpenStudy (anonymous):

oh i forgot the x

OpenStudy (anonymous):

im realy stuck with this one solve for x \[-2x-1=-11\] \[4x \over 5\] =8

OpenStudy (anonymous):

-2x-1=-11 move -2x to the right handside it becomes 2x -1=-11+2x move -11 to the left handside it becomes 11 11-1=2x 10=2x divide both sides by 2 5=x

OpenStudy (anonymous):

thank you so much hope to chat again

OpenStudy (anonymous):

nope